OpenStudy (anonymous):

Polar Company has designed an ice block in the shape of a cube. The volume of the cube is 15 625 cm3. Which of the following dimensions is the smallest opening of an ice dispenser that will accommodate length AB?

OpenStudy (anonymous):

25cm

OpenStudy (anonymous):

\[\sqrt[3]{15625}=25\]

OpenStudy (anonymous):



OpenStudy (anonymous):

bottom corner to the top corner at the opposite end is AB

OpenStudy (anonymous):

Therefore opening needs to be 25x25 (assuming it will always fall base first. If not look at the image above)

OpenStudy (anonymous):

in which case \[25\sqrt{2}\times25\sqrt{2}\]